OpenStudy (1dfanatic):

The length of a rectangle is 5 in. more than its width. The perimeter of the rectangle is 42 in. A. Define a variable for the width. Write an expression for the length in terms of the width. B. Use the formula for the perimeter of a rectangle to write an equation to find the width of the rectangle. Solve your equation for the width. C. What is the length of the rectangle?

OpenStudy (anonymous):

A.) Let width = w and length = l. w + 5 = l

OpenStudy (anonymous):

B.) P = 2l + 2w. 42 = 2(w + 5) + 2w

OpenStudy (anonymous):

Are you able to solve from there?